OpenStudy (yessenia27):

What is the value of a in the equation a = 2 + 3a + 10?

rebeccaxhawaii (rebeccaxhawaii):

move like terms

rebeccaxhawaii (rebeccaxhawaii):

a-3a=2+10

alones (alones):

yes yes move like terms

rebeccaxhawaii (rebeccaxhawaii):

whats a-3a=? whats 2+10=?

OpenStudy (yessenia27):

Huh

rebeccaxhawaii (rebeccaxhawaii):

what is a-3a or 1a-3a=?

alones (alones):

simplify both sides of the equation a=2+3a+10 a=3a+12 than subtract 3a from both sides. a−3a=3a+12−3a −2a=12 and divide both sides by -2. \( \frac{ -2a }{ -2 }\) =\(\frac{ 12 }{ -2 }\) a=−6